# Simplified functional flow equation

### v2

@@ Abstract
We adapt the precise definition of the flowing effective action in order to obtain a functional flow equation with simple properties close to physical intuition. The simplified flow equation is invariant under local gauge transformations and suitable for both euclidean and Minkowski signature and analytic continuation. The cutoff always removes fluctuations close to zeros of the inverse full propagator. A formulation of the simplified flow equation in terms of renormalized scale invariant fields permits direct access to scaling solutions and associated fixed points. Corrections to the simplified flow equation involve a field-dependent modification of the cutoff for which we discuss a systematic expansion. Truncated solutions for a scalar field theory in four dimensions suggest a new fixed point with a field-dependent coefficient of the kinetic term.
@@
We propose an optimization of the definition of the flowing action for which the simplified flow equation can be considered as the starting point of a systematic expansion. We compute explicitly the general form of the corrections to the simplified flow equation. For our definition of the flowing effective action the exact flow equation can be formulated in an implicit form. The expansion around the simplified flow equation has the character of a type of loop expansion for a correction term. The simplified flow equation itself is non-perturbative. It includes already all one-loop effects exactly. Having a structure similar to the exact flow equation for the effective average action the simplified flow equation also sums up higher loops and can be used for a full non-perturbative analysis. The loop expansion only concerns corrections to the precise shape of the infrared cutoff.
Optimization by the choice of a cutoff term which is quadratic in the fluctuation field and independent of the macroscopic field has already a long tradition [23]. We aim here for a more general optimization for which the cutoff term is allowed to depend on the macroscopic fields [24, 25]. These macroscopic fields are the expectation values of the microscopic fields over which one integrates in the functional integral. The price to pay is that the one-loop form of the exact flow equation is no longer guaranteed to hold. We propose an optimized flow equation which maintains the simple one-loop form. The propagator in the loop involves, however, field-dependent corrections to the effective infrared cutoff which are, in turn, given by a loop expansion.
@@
For a single real scalar field $\varphi(x)$ the flowing action $\Gamma_{k}[\varphi]$ is a functional of this “macroscopic field”. It flows in dependence on the renormalization scale $k$. In the simplest approximation in flat Euclidean space the simplified functional flow equation reads
For $d$ even dimensions a typical value is $n=d/2+1$. With $\Gamma_{k}^{(2)}$ the matrix of second functional derivatives of $\Gamma_{k}$ the r.h.s. of eq. (1) involves the $n$-th power of the regularized full field-dependent propagator. The quantity $Z$ denotes a suitable wave function renormalization, with anomalous dimension $\eta=-k\partial_{k}\ln Z$. Below we will generalize the wave function renormalization $Z$ to a field-dependent quantity. This replaces $Z^{-1}\Gamma_{k}^{(2)}$ by a renormalized version of the second functional derivative, which is an important characteristic of our setting.
Eq. (1) has been proposed in [26] for the Wilsonian action which differs conceptually from the flowing effective action. In this context it follows from a version of heat kernel regularization. Even though heat-kernel evaluations of the simplified flow equation are possible, we do not base our approach on a heat kernel regularization.
@@
We next introduce the general form of the simplified flow equation. For several fields $\varphi$ becomes a vector in field space and $\Gamma_{k}^{(2)}$ a corresponding matrix. The simplified flow equation is defined by
@@
The matrix $N[\varphi]$ generalizes $Z^{1/2}$ to a type of field-dependent wave function renormalization. Correspondingly, the matrix $E[\varphi]$ generalizes $\eta$ in eq. (1). We take
@@
such that for a $k$-dependence of $NN^{T}$ of the form $N[\varphi]N^{T}[\varphi]=A[\varphi]k^{-\eta}$ one indeed finds $E[\varphi]=\eta$. This includes the field-independent case $NN^{T}=Z$. For an arbitrary metric $g_{\mu\nu}$, with $g=\det(g_{\mu\nu})$, the factor $g^{-1/2}$ takes out the factor $g^{1/2}$ arising from $\Gamma=\int_{x}\sqrt{g}L$. In flat Euclidean space one has $g_{\mu\nu}=\delta_{\mu\nu}$ and $g=1$, while for Minkowski space $g_{\mu\nu}=\eta_{\mu\nu}$ implies $g=-1$. The factor $\sqrt{g}$ takes care of factors of $i$. The optimized flow equation (6) can therefore be used for an arbitrary signature of the metric. Analytic continuation can be performed by analytic continuation of the metric, keeping the coordinates fixed. All factors of $i$ are accounted for by the metric.
@@
Finally, the flow equation (6) offers an important technical advantage:
@@
For a simple enough form of $\Gamma_{R}^{(2)}$, for example in a derivative expansion, one can use standard heat kernel expressions for the trace in eq. (10). An example is the computation of the flow of a field-dependent effective Planck mass in quantum gravity [27]. For gauge invariant $\Gamma_{k}$ the second functional derivative is gauge covariant, which makes the gauge invariance of $B(z)$ and therefore the flow generator manifest. A cutoff depending on the second functional derivative of an effective action and its relation to heat kernels has been discussed for the background field formalism in ref. [28, 29]. For other proposals of non-perturbative flow equations using heat kernels see refs. [30, 31, 32, 33, 34, 35, 26, 36].
As a second technical advantage we may state
  1. (vii)
@@
This flow equation would be exact only if $R_{k}$ is independent of $\varphi$. In general, this is not the case here since both matrices $N$ and $B$ are allowed to depend on $\varphi$. The corrections to eq. (11) for field-dependent $B$ and $N$ will be addressed below. We may consider eq. (11) as the flow equation in the background formalism for which the correction term discussed in ref. [8] has been omitted. Alternatively, it corresponds to the gauge invariant flow equation without the correction terms specified in refs. [24, 25].
@@
In summary, the simple analytic structure of the simplified flow equation (6) is achieved by the adaption of the cutoff function. We see no obvious reason why similar solutions of the differential equation (18) should not exist for $\gamma\neq 1$. This possibly extends to the inclusion of corrections to $\tilde{r}$. For an optimal choice of the cutoff function we believe that the dominant corrections to the simplified flow equation arises from the corrections to eq. (11) for field-dependent $R_{k}$. They add to $R_{k}$ in eq. (11) a correction term $Q_{k}$, as we will establish below in eq. (61).
@@
in accordance with eq. (Simplified functional flow equation).
The sum runs over all particles with renormalized dimensionless mass term $\tilde{m}_{i}^{2}=m_{i}^{2}/k^{2}$. Only particles with small enough $\tilde{m}_{i}^{2}$ contribute to the flow in a given range of $k$. For the scalar fluctuations there are three Goldstone modes with $\tilde{m}_{i}^{2}=u^{\prime}$, while the radial mode obeys $\tilde{m}^{2}=u^{\prime}+2\tilde{\rho}u^{\prime\prime}$. Here primes denote derivatives with respect to $\tilde{\rho}$. For the contribution from fermion fluctuations one has an additional minus sign and $\tilde{m}_{i}^{2}=y_{i}^{2}\tilde{\rho}$, with $y_{i}$ the Yukawa couplings. (Each Majorana fermion contributes a factor $-2$ such that for a Dirac fermion there is a factor $-4$.) For the two $W$-Bosons one finds $\tilde{m}_{i}^{2}=g_{2}^{2}\tilde{\rho}/2$ and for the $Z$-boson $\tilde{m}_{i}^{2}=g_{2}^{2}\tilde{\rho}/2+3g_{1}\tilde{\rho}/10$, with $g_{2}^{2}$ and $g_{1}^{2}$ the gauge couplings of the electroweak gauge groups $SU(2)$ and $U(1)$. For each gauge boson there is a factor three corresponding to the number of physical degrees of freedom for a massive gauge boson. The measure factor subtracts for each gauge boson the contribution of one field with $\tilde{m}_{i}^{2}=0$. At the minimum of the potential this cancels the contribution of the scalar Goldstone modes, such that only contributions from physical fluctuations remain. The contribution from physical metric fluctuations can be included in a similar way [37, 38, 39]. Eq. (Simplified functional flow equation) holds for an arbitrary constant metric, and therefore both for Euclidean or Minkowski flat space.
@@
For homogeneous $m^{2}$ we recover the flow of the potential in eq. (Simplified functional flow equation).

### v3

@@ Abstract
We adapt the precise definition of the flowing effective action in order to obtain a functional flow equation with simple properties close to physical intuition. The simplified flow equation is invariant under local gauge transformations and suitable for both euclidean and Minkowski signature and analytic continuation. The cutoff always removes fluctuations close to zeros of the inverse full propagator. A formulation of the simplified flow equation in terms of renormalized scale invariant fields permits direct access to scaling solutions and associated fixed points. Our setting is based on a particular choice of cutoff function which depends on the macroscopic fields. Corrections to the simplified flow equation involve a field-dependent modification of the cutoff for which we discuss a systematic expansion. Truncated solutions for a scalar field theory in four dimensions suggest a new fixed point with a field-dependent coefficient of the kinetic term.
@@
We propose an optimization of the definition of the flowing action for which the simplified flow equation can be considered as the starting point of a systematic expansion. The key ingredient is the dependence of the cutoff function on the macroscopic fields. This induces corrections to the flow equation beyond the one loop form of ref. [7]. We modify the source term in order to cast the corrections into a simple form. We compute explicitly the general form of the corrections to the simplified flow equation. For our definition of the flowing effective action the exact flow equation can be formulated in an implicit form. The expansion around the simplified flow equation has the character of a type of loop expansion for a correction term. The simplified flow equation itself is non-perturbative. It includes already all one-loop effects exactly. Having a structure similar to the exact flow equation for the effective average action the simplified flow equation also sums up higher perturbative loops and can be used for a full non-perturbative analysis. The loop expansion proposed here only concerns corrections to the precise shape of the infrared cutoff.
Optimization by the choice of a cutoff term which is quadratic in the fluctuation field and independent of the macroscopic field has already a long tradition [23]. We aim here for a more general optimization for which the cutoff term is allowed to depend on the macroscopic fields [24, 25]. These macroscopic fields are the expectation values of the microscopic fields over which one integrates in the functional integral. The price to pay is that the one-loop form of the exact flow equation is no longer guaranteed to hold. We propose an optimized flow equation which maintains formally the simple one-loop form. The propagator in the loop involves, however, field-dependent corrections to the effective infrared cutoff which are, in turn, given by a loop expansion.
@@
The simplified flow equation is defined by
@@
Here $\varphi$ is the macroscopic field, which is a vector in the space of internal or Lorentz indices, and $\Gamma_{k}^{(2)}[\varphi]$ denotes the second functional derivative of the flowing action with respect to $\varphi$. The matrix $N[\varphi]$is a type of field-dependent wave function renormalization. Correspondingly, the matrix $E[\varphi]$ is a field-dependent anomalous dimension. We take
@@
such that for a $k$-dependence of $NN^{T}$ of the form $N[\varphi]N^{T}[\varphi]=A[\varphi]k^{-\eta}$ one finds $E[\varphi]=\eta$. The field-dependent factors $N[\varphi]$ become important if the kinetic term in $\Gamma_{k}$ is multiplied by a field-dependent function $\sim NN^{T}$, as for a field-dependent Planck mass in quantum gravity, or the field-dependent “kinetial” for the suggested new fixed point in scalar quantum field theory. The trace in eq. (1) amounts to a momentum integration or an integration over positions. We take $n=3$ for $d=3,4$ and $n=2$ for $d=1,2$, such that the momentum integral is finite for $\Gamma_{k}^{(2)}\sim q^{2}$.
For an arbitrary metric $g_{\mu\nu}$, with $g=\det(g_{\mu\nu})$, the factor $g^{-1/2}$ takes out the factor $g^{1/2}$ arising from $\Gamma=\int_{x}\sqrt{g}L$. In flat Euclidean space one has $g_{\mu\nu}=\delta_{\mu\nu}$ and $g=1$, while for Minkowski space $g_{\mu\nu}=\eta_{\mu\nu}$ implies $g=-1$. The factor $\sqrt{g}$ takes care of factors of $i$. The optimized flow equation (1) can therefore be used for an arbitrary signature of the metric. Analytic continuation can be performed by analytic continuation of the metric, keeping the coordinates fixed. All factors of $i$ are accounted for by the metric.
@@
As as technical advantage we may state
@@
Finally, the flow equation (1) offers one more important technical advantage:
  1. (vii)
@@
For a simple enough form of $\Gamma_{R}^{(2)}$, for example in a derivative expansion, one can use standard heat kernel expressions for the trace in eq. (5). An example is the computation of the flow of a field-dependent effective Planck mass in quantum gravity [26]. For gauge invariant $\Gamma_{k}$ the second functional derivative is gauge covariant, which makes the gauge invariance of $B(z)$ and therefore the flow generator manifest.
An interesting approximation to the simplified flow equation considers a simple scalar field $\varphi$ and approximates $NN^{T}$ by a field-independent wave function renormalization $Z$. For this approximation in flat Euclidean space the simplified functional flow equation reads
This equation has been obtained before from the proper time flow equation [27, 28, 29, 30, 28, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40]. It has also been proposed in [41] for the Wilsonian action, which differs conceptually from the flowing effective action.
Our setting is based on a functional integral for the flowing action and we do not use a regularization of heat kernels for the definition of the flowing action. The precise connection which leads to the coincidence of eq. (6) with a proper time flow equation is not known to us. This may constitute an interesting topic of research. One also may raise the question if there exist possible connections between the full simplified flow equation (1) and the corrections to it with a regularization of heat kernels.
The simplified flow equation can also be compared with the background formalism for the effective average action [8]. The background formalism involves in addition to the macroscopic field $\varphi$ a background field $\bar{\varphi}$ on which the cutoff may depend. The flowing action $\Gamma_{k}[\varphi,\bar{\varphi}]$ is then a functional of two fields, for example two gauge fields $A_{\mu}$ and $\bar{A}_{\mu}$. Identifying $\bar{A}=A$ the effective action $\Gamma_{k}[A]=\Gamma_{k}[A,A]$ is gauge invariant if the cutoff function involves covariant derivatives with respect to the background field $\bar{A}$. The second functional derivative $\Gamma_{k}^{(2)}[A,\bar{A}]$ which appears in the exact one loop form of the flow equation is taken at fixed $\bar{A}$. A commonly used approximation replaces $\Gamma_{k}^{(2)}[A,\bar{A}]$ by $\Gamma_{k}^{(2)}[A]$. This yields a flow equation with a similar structure as the simplified flow equation. In fact, the background formalism equals the flowing action discussed in this note up to a modification of the source term. The structure of the corrections to the exact flow equation in the background field formalism is known [8]. They are rather complicated since two gauge fields $A$ and $\bar{A}$ are involved. The corrections to the simplified flow equation have a simpler structure. A cutoff depending on the second functional derivative of an effective action and its relation to heat kernels has been discussed for the background field formalism in ref. [42, 43].
@@
This flow equation would be exact only if $R_{k}$ is independent of $\varphi$. In general, this is not the case here since both matrices $N$ and $B$ are allowed to depend on $\varphi$. The corrections to eq. (11) for field-dependent $B$ and $N$ will be addressed below. We may also consider eq. (11) as the flow equation in the background formalism for which the correction term discussed in ref. [8] has been omitted. Alternatively, it corresponds to the gauge invariant flow equation without the correction terms specified in refs. [24, 25].
@@
In summary, the simple analytic structure of the simplified flow equation (1) is achieved by the adaption of the cutoff function. We see no obvious reason why similar solutions of the differential equation (18) should not exist for $\gamma\neq 1$. This possibly extends to the inclusion of corrections to eq. (13) or more generally to $r$. For an optimal choice of the cutoff function we believe that the dominant corrections to the simplified flow equation arises from the corrections to eq. (11) for field-dependent $R_{k}$. They add to $R_{k}$ in eq. (11) a correction term $Q_{k}$, as we will establish below in eq. (61).
@@
in accordance with eq. (Simplified flow equation).
The sum runs over all particles with renormalized dimensionless mass term $\tilde{m}_{i}^{2}=m_{i}^{2}/k^{2}$. Only particles with small enough $\tilde{m}_{i}^{2}$ contribute to the flow in a given range of $k$. For the scalar fluctuations there are three Goldstone modes with $\tilde{m}_{i}^{2}=u^{\prime}$, while the radial mode obeys $\tilde{m}^{2}=u^{\prime}+2\tilde{\rho}u^{\prime\prime}$. Here primes denote derivatives with respect to $\tilde{\rho}$. For the contribution from fermion fluctuations one has an additional minus sign and $\tilde{m}_{i}^{2}=y_{i}^{2}\tilde{\rho}$, with $y_{i}$ the Yukawa couplings. (Each Majorana fermion contributes a factor $-2$ such that for a Dirac fermion there is a factor $-4$.) For the two $W$-Bosons one finds $\tilde{m}_{i}^{2}=g_{2}^{2}\tilde{\rho}/2$ and for the $Z$-boson $\tilde{m}_{i}^{2}=g_{2}^{2}\tilde{\rho}/2+3g_{1}\tilde{\rho}/10$, with $g_{2}^{2}$ and $g_{1}^{2}$ the gauge couplings of the electroweak gauge groups $SU(2)$ and $U(1)$. For each gauge boson there is a factor three corresponding to the number of physical degrees of freedom for a massive gauge boson. The measure factor subtracts for each gauge boson the contribution of one field with $\tilde{m}_{i}^{2}=0$. At the minimum of the potential this cancels the contribution of the scalar Goldstone modes, such that only contributions from physical fluctuations remain. The contribution from physical metric fluctuations can be included in a similar way [44, 45, 46]. Eq. (Simplified flow equation) holds for an arbitrary constant metric, and therefore both for Euclidean or Minkowski flat space.
@@
For homogeneous $m^{2}$ we recover the flow of the potential in eq. (Simplified flow equation).